OpenStudy (anonymous):

Pete, a professional bowler, is unhappy with any game below 200. Over time, 80% of his games exceed this score. What is the probability that Pete exceeds 200 in at least 9 of his next 10 games?

OpenStudy (anonymous):

I have: P(9/10 games> 200) +P(10/10 grames> 200) Or 1-P(Doesn't exceed any or doesn't exceed 1 game)

OpenStudy (anonymous):

1st way is easiest on this one.

OpenStudy (anonymous):

Do I use combinatorial

OpenStudy (anonymous):

if you do the 1 - P thing you'll have to calulate P(8/10 >200) P( 7/10>200), ... all the way down to P(0/10 >200)

OpenStudy (anonymous):

at least 9 means nine or ten easier than computing none one two three four five six seven eight

OpenStudy (anonymous):

use the binomial to compute this but the only combination you need is ten choose 9 which is ten

OpenStudy (anonymous):

Start with the easier one. What's P(10/10 games are >200)?

OpenStudy (anonymous):

.8^10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

right. and P(9/10 games are >200) = ? Any idea there?

OpenStudy (anonymous):

.8^9 *.2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

almost. take what you have there and multiply by 10. There are 10 possibilities for the game that you are below 200. The 1st game, 2nd game, ... , 10th game.

OpenStudy (anonymous):

.8^9 * 2?

OpenStudy (anonymous):

10 * .8^9 * 2

OpenStudy (anonymous):

Final answer: .8^10 + 10 * .8^9 * 2 = whatever that is

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

means of course \[10\times (.8)^9\times .2\]